A monoatomic gas undergoes a process in which the ratio of P to V at any instant is constant and equals to 1. What is the molar heat capacity of the gas??

Options:
A)4R/2 B)3R/2 C)5r/2 D)0

Grade:12

4 Answers

mohit yadav
14 years ago

HI

I THINK DA ANSWER 2 DIS QUESTION IZ [A]4R/2

DIS IZ AN ISOTROPIC PROCESS (defiined as process in which PV*=constant hereV*meansV ki power x)AND FORMULA OF MOLAR HEAT CAPACITY FOR SUCH A PROCESS IZ=R/y-1-R/x-1 y=5/3 for a monoatomic ideal gas and here P/V=CONSTANT SO x=-1 so ANSWER=3R/2-(-R/2)=4R/2. U CAN FIND DERIVATION OF DIS FORMULA IN HCV . PLZ REPLY ME IF D ANSWER IS CORRECT OR NOT.

ShipWreck GrimJaw
14 years ago

Thanks very much Mohit. The answer is A) 4R/2. But could you please tell me what you exactly meant by HCV and where exactly to look.

 

 

                                          

mohit yadav
14 years ago

HI

HCV means Physics book of HC VERMA ,refer dere 2 get dis formula ,available may b in ur school library or buy this book as dis iz d best book 2 refer for IIT-JEE  refer 2 volume-2 of HCV.U will surely enjoy reading dat book .U may also refer to some physical chemistry book.

BEST OF LUCK 2 U FOR UR ENTRANCE EXAMS.

Brijesh Rohit
5 years ago
For a polytropic process
Cm = Cvm + R/(1-x) 
where x is the polytropic index. Given that
P/V =1 hence x=-1,
Therefore Cm= 3R/2 + R/[1-(1)]  = 2R.